OpenStudy (anonymous):

4) Solve: -7x > 14 * x > -2 x < 2 x > -28 x < -2 @StudyGurl14 @EBOLA can you help

OpenStudy (anonymous):

its a also medal me

OpenStudy (anonymous):

yep it is

OpenStudy (studygurl14):

Divide both side by -7, then flip the sign because you are dividing by a negative number

OpenStudy (anonymous):

ok so it would be x>-2 @StudyGurl14

OpenStudy (studygurl14):

not quite. flip the sign... x < -2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anyone need help just tag me

OpenStudy (anonymous):

can you help with more

OpenStudy (anonymous):

who me??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7) Admission to The Arcade Room includes a bonus card with 100 points that can be spent on games. Games cost 6 points each. Which of the following equations shows the relationship between the number of games and the points left on the bonus card? (p stands for points and g stands for games) * p = 100 + 6g p = 100 - 6g 100 - 6 = p 6 + g = p

OpenStudy (anonymous):

i think this one is p=100-6g

OpenStudy (anonymous):

NO!! im not help you leave me alone!!!

OpenStudy (anonymous):

im joking

OpenStudy (anonymous):

:)

OpenStudy (studygurl14):

You're correct @cool41 Good job

OpenStudy (anonymous):

you're right good job

OpenStudy (anonymous):

thanks @StudyGurl14

OpenStudy (studygurl14):

:)

OpenStudy (anonymous):

and @abswag13

OpenStudy (anonymous):

more

OpenStudy (anonymous):

8) Simplify: 7x + 4y - 3x + 10z + 5y * 4x + 9y + 10z 4x + 19y 23xyz 10x + 10z + 9y

OpenStudy (anonymous):

i think it is 4x+9y+10z?

OpenStudy (studygurl14):

correct :)

OpenStudy (anonymous):

yayy thanks thats all for now have to go to cxlass

TheSmartOne (thesmartone):

Also 1 question per post.. Because what if 3 different people helped you with your 3 questions. You can't give them all medals.

OpenStudy (studygurl14):

I know ^ @TheSmartOne I hate when people do that...but it is so hard to get them to do that

TheSmartOne (thesmartone):

:)